OpenStudy (anonymous):

x intercept of y = -x^2 + 4

OpenStudy (dumbcow):

replace y with 0 x^2 = 4

OpenStudy (anonymous):

y=-x^2+4=4-x^2=(2-x)(2+x), so x intercepts are x=2 or x=-2